OpenStudy (anonymous):

Are 0/infinity, infinity/0, -infinity/+infinity, -infinity^0, 0*-infinity, intermediate form types? @Mathematics Are 0/infinity, infinity/0, -infinity/+infinity, -infinity^0, 0*-infinity, intermediate form types? @Mathematics @Calculus1

OpenStudy (anonymous):

\[\frac{0}{\infty}\] is just 0

OpenStudy (anonymous):

so no it is not "indeterminate"

OpenStudy (anonymous):

likewise neither is \[\frac{\infty}{0}\]

OpenStudy (anonymous):

the rest are "indeterminate"

OpenStudy (anonymous):

Thanks :D

OpenStudy (anonymous):

yw